OpenStudy (anonymous):

what is the antiderivate of 9cos3t???

Parth (parthkohli):

integrate it

Parth (parthkohli):

In other words, what is \(9\cos3t\) the derivative of?

Parth (parthkohli):

of course with respect to \(t\).

OpenStudy (anonymous):

Hint: \[\int\limits_{}^{} \cos ax dx = \frac{ \sin ax }{ a } + c\]

Parth (parthkohli):

You start by taking the constant out.\[9\int\cos (3t)\]Use what Denebal gave.

Parth (parthkohli):

Denebel*

Parth (parthkohli):

By the way, you can also take \(u = 3t\) :)

OpenStudy (anonymous):

would it be sin3/2t^2

Parth (parthkohli):

Erm, no.

Parth (parthkohli):

\[9\int \cos(3t)dt\]Sorry for the mistake above

OpenStudy (anonymous):

Where do you get 2t^2 ???

OpenStudy (anonymous):

the antiderivate of 3t is 3/2t^2

Parth (parthkohli):

\[9 \times{1 \over 3}\int \cos(u)du\]\[3\int \cos(u)du\]\[\implies 3\times \sin(u) + C\]

OpenStudy (anonymous):

\[9\int\limits_{}^{} \cos (3t) dt = \frac{ 9\sin(3t) }{ 3 } = 3\sin(3t) + C\]

Parth (parthkohli):

Substituting \(u\) back,\[3\sin(3t) + C\]Yep

OpenStudy (anonymous):

thank you!... i can't do antiderivativs when they have a sin, cos, etc in them

OpenStudy (anonymous):

∫ cosu dx = sinu / u' Does it make any sense to you?

OpenStudy (anonymous):

or ∫cos ax dx = sin ax / a + C Does it make any sense at all?

OpenStudy (anonymous):

yeah, thank you